OpenStudy (anonymous):

What are the domain and range of the function y=2 sqrt(3x +4) -5?

OpenStudy (anonymous):

domain is possible values of x we can take in square root the number must be >= 0 always so 3x+4>=0 3x>=-4 x>=-4/3 (domain) range is the output we get

OpenStudy (anonymous):

\[y=2*\sqrt{3x+4}-5\]

OpenStudy (anonymous):

its min value is when 3x+4=0 that is y=2*0-5=-5 range>= -5

OpenStudy (anonymous):

@jerryK

OpenStudy (anonymous):

Yeah, @g17 has it right..